A1459 3 min on/offs with lots of baseline ripple.
(new Airconditioner causing problems)
07sep01. The baseline ripples
(vs freq) and the temperature variation of the room.
08sep01. Turret room temperature variation.
08sep01. Check cband,
lband narrow, lband wide for jumps in total power.
14sep01. Total power
jumps coming from new airconditioner compressor.
20sep01. lbn Image while
turning air conditioner on/off.
28sep01. lbn image with magnetic
contactor installed on new ac compressor.
05oct01. Turn louvre motor, blower,
Ac compressor, 208Volt breaker on/off.
31oct01. Timer module for compressor causing the
A1459. The baseline
A1459 is a search for galaxies in the zone of avoidance
using lband narrow covering 1350-1420 Mhz. 3 minute position switching
was used. The experiment ran from 18. to 23:30 hours AST starting on 30aug01.
They had baselines with lots of ripples in them. The figures show their
results for the days 30aug01 through 07sep01 (the experiment continues
on after this but i only analyzed the first 9 days)
plots (5mb ps file) and polB
plots (5mb ps file) show on/off-1 in kelvins for each day. For each
day there are 4 frequencies. Each strip has been offset from the previous
by .2 kelvins for display purposes. The baseline ripple is:
Bad means that there was ripple in a large fraction of the observations.
Some of the better days had large ripple in a few strips. The ripple was
on the order of the strip offset (.2 Kelvins). The 25 Mhz integration
over 180 seconds with hanning smoothing (and no addition of polarizations)
should have given 30K*sqrt(2)/sqrt(25e6/1024*2*180)=.015 Kelvins. Some
properties of the ripples were:
30aug, and 04sep->07sep were worse than 31aug->03sep.
The ripples are in all 4 sbc so it is probably in or before the 1st IF
(or in the correlator).
It is on both polarizations so it is not a single amplifier/mixer in the
The ripples in polB are shifted about 2.5 Mhz down from the same ripple
in polA. Look at 07sep01 1396.4 polA (6th strip from the top) and 1394
polB (6th strip from the top).
We took data at cband tracking blank sky. We took adjacent 3 minute sections
and computed on/off-1. 10 of these on/off-1 are shown in the
plot (each offset for plotting). The vertical scale is Tsys rather
than Kelvins. There is a 1 Mhz standing wave (since we did not do position
switching) but the rms is < .001*( tsys=30) <. 03 Kelvins. This is
close to the theoretical value and an order of magnitude less than the
bad ripples at lband. This rules out the correlator as the problem. The
setup differences were the 1st IF (lb 750Mhz if, cband 1.5 Ghz IF), dewar/postamps,
The ripples are across all of the 100 Mhz. If it is rfi , it must be outside
the IF bandpass and driving something into saturation. It could be a noise
diode from one of the other receivers that is firing when it isn't supposed
to (we had this problem originally when the ttl cal triggered all of the
receivers at the same time way back in 199?).
Turret room temperature variation. (topa1459)
There is a temperature sensor in the gregorian turret
room. It measures the air temperature in the middle of the room. The temperature
plots 30aug01 -> 07sep01 19:00 to 24:00 show the room temperature
(there is no data for 05sep01). There is a temperature cycle of about 10
minutes for the "bad" days while the good days have a temperature drift.
The 10 minute cycle has an amplitude of 1 deg F (air temperature). On 09sep01
both air conditioners were turned off and the temperature rose from 78
to 95 degrees of a few hours. During this time, the baselines had no ripples.
So there is a definite correlation between the air conditioner cycling
and the baseline ripples. It is either temperature or some indirect cause
like (e.g.. running the air conditioner could be changing the line
voltage that the if/lo sees).
A new air conditioner was added to the turret room
in early aug01. There are two of them set at different thresholds. The
plot shows the room
temperature 01aug01 (bottom) through 08sep01 (top) (2mb psfile). Each
day has been offset by 2 degF for plotting. The new second air conditioner
was installed around 8aug01. With the new air conditioner, the temperature
cycle is now about 1 degree over 10 minutes.
Total power variation
of cband and lband narrow with temperature. (topa1459)
On 7sep01 and 08sep01 blank sky was tracked using the
cband, lband narrow, and lband wide systems. Data was sampled at 1 second
intervals with the correlator. The goal was to check the system stability
after the reported baseline problems.
figures show total power time series sampled at once a second. There
is a trace for each sbc (freq) and polarization.
Lbn has discrete jumps in power correlated with temp. Cband does not. Lband
wide may or may not have the jumps. The frequency spectrum of the jumps
show .5 Mhz stripes.
The cband data showed no jumps in power.
Figure 1 top has the lband narrow total power samples taken on 08sep01.
The lower plot is a blowup after averaging all 4 sub correlators (freq)
for each polarization. Near the bottom of each temperature cycle there
is a jump of about .3% Tsys down by all sub correlators and both polarizations.
About a third of the way up from the bottom of each temperature cycle there
is a corresponding jump upwards of about .3% Tsys. With a Tsys of
28K this is about .1 Kelvins. Since there are two air conditioners in the
room, it looks like one of the air conditioners is turning off causing
the downward jump. The temperature rises, the air conditioner turns back
on causing a positive jump in the total power. The two air conditioners
have their turn off thresholds set to 73.4F (23 C) and 71.6f (22 C). Their
turn on value is probably a fixed value above this. The temperatures differ
from those plotted since the room temperature sensor is located in the
center of the room while the air conditioner sensor is measured at the
air intake. On top of the discreet jumps, there is the same temperature
variation of gain as cband.
Figure 2 is a blowup of 800 seconds of figure 2 data 72800 through 73600
seconds from midnight. Plotted is sbc 1, polA (total power-median(totalpower))*500.
The temperature is over plotted in color. An
image was made (200kb gif) of the 800 spectra that correspond
to these total power points. The spectra was bandpass corrected using the
average bandpass. The vertical axis on the image (seconds from midnight)
corresponds to the horizontal access of the line plot. Looking at
the image you can see that the jumps are not constant in frequency.
There are .5 mb wide stripes distributed across the band.
Figure 3 has lband wide total power samples. The receiver had the hybrid
out and only pol A is plotted (pol B has a large 100 second oscillation).
The 1% per degF power fluctuation is also present in this receiver.
You can see the 100 sec oscillation in this polarization as well. There
may also be a .3% jump when the temperature cycle goes through the minimum
but it is difficult to say because of the other junk going on.
14sep01 total power jumps
while air conditioner power cycled. (topa1459)
On 14sep01 blank sky was tracked during
the day with the lband narrow receiver. A 25Mhz bandwidth was sampled at
a 1 hz rate close to 1418 Mhz. While the data was being taken the two airconditioning
units had there power turned on and off (at different times) to see how
they affected the total power.
figure shows the results. It plots the total power versus time. The
colored vertical lines are:
vertical solid red. new airconditioner turned on.
vertical dashed red. old airconditioner turned on.
vertical solid green. new airconditioner turned off.
vertical dashed green. old airconditioner turned off.
These are split level units with the consoles in the turret room and
the compressors on the service platform. The power to the consoles was
turned on/off with a remote unit. Between 53 and 80 seconds after the new
Ac was turned on, the rf power jumped by .3%. This did not occur with the
older unit. The new unit would then turn it's compressor off before we
shut off the power. It turns out that the evaporator unit was icing up
so it was not running continuously (even though the temperature in the
room was 90 F).
The jumps in the lband narrow receiver are caused by the new Ac compressor
20sep01 lbn image of air conditioning
On 20sep01 during the day the lband narrow receiver was used to take
data while the new airconditioner unit was turned on and off. 1024 channels
were taken over a 25 Mhz bw at 1 second integrations. There were 1060 seconds
of data. The image was flattened in frequency by using the time average
of the data. It was flattend in time by using 50 channels away from the
image shows on/off-1 versus frequency and time. There are 3 sets of
dashed lines added to the plot:
Using 1417 Mhz as a reference you can see that:
At 36 seconds the console was turned on.
At 50 seconds the compressor was heard to come on.
At 100 seconds the console/compressor turned off. The image becomes dark
implying that the power has gone done.
At 150 seconds the console was turned on. No rfi yet.
At 190 seconds the rfi appeared but they did not hear the compressor turn
At 400 seconds the console/compressor was turned off and the rfi went away.
At 450 seconds the console was turned on. No rfi yet.
At 590 seconds the compressor was heard to come on and the rfi appeared.
At 750 seconds the console/compressor was turned off and the rfi
At 840 seconds the console was turned on. No rfi yet.
At 940 seconds the rfi appeared but they did not hear the compressor start.
At 1060 they heard the compressor come on.
short dashes. The remote control was used to turn on the airconditioner
midsized dashes. A person watching the compressor heard it come on.
long dashes. The remote control was used to turn off the console and compresssor.
So the rfi is always turning off with the console/compressor. Twice
the rfi started when the compressor started. Two times the rfi started
without hearing the compressor turn on. The other air conditioner was running
at the time. I doubt that the rfi is fromthe old air conditioner since
it was not turned off when the new unit was (and the rfi always disappeared).
28sep01 New magnetic contactor
on ac installed. (topa1459)
The airconditioner compressor came with a magnetic contactor to start
the compressor. This contactor was defective on the new ac compressor
and it was replaced with a solid state (SCR) switch. The SCR switch could
have been generating noise so on 28sep01 it was replaced with another
magnetic contactor. The image
shows lbn tracking blank sky. The image processing and embedded horizontal
lines are the same as the 20sep01 image above. For this experiment the
old air conditioner was shut off.
Small dots are console on.
The next largest set of dashes is compressor on
The next largest set (8 dashes/page) is everything off.
The area between the largest dashes (2 dashes per page) was used for the
There are ripples that start right after the console turns on and go
off when the compressor goes off so it is not just the compressor. The
area 800-1000 seconds looks the flattest since it was used for the
bandpass correction. Since we were tracking blank sky, as we move away
from this location the normal standing waves increase. This image in idl
shows the start/stop at the console on/off a lot more clearly.
shows total power versus time for the above image (and two other
frequency bands). Three different 25 Mhz bands are over plotted (1420,1400,1380).
The vertical scale is totalPower[i]/median(totalpower) - 1.
The vertical colored lines are:
All three separate bands jumped up 10 seconds after the console came on.
They all jumped back down when the console/compressor went off. The
compressor did not come on till much later so the problem is not in the
Red - air conditioner console on (small dots in image).
Green- air conditioner compressor comes on (medium dashes in image).
Blue - console/compressor off (8 dashes per page in image).
The next thing to check is the louvre motor that causes the louvres
to swing back and forth and the blower motor. Both of these turn on shortly
after the console is turned on.
05oct01 New air conditioner, louvre motor,
ac breaker on/off. (topa1459)
Blank sky was tracked with lbn. 25 Mhz bandpasses were dumped once
a second centered at 1420,1400,1380 Mhz. 5 separate 300 second integrations
were performed. The old air conditioner was turned off. After the first
plot the breaker for the old air conditioner was also shut off. Figure
1 shows the total power versus time (normalized to the median total
power) for each of the 5 observations. A 1st order polynomial removed the
za dependence. Color vertical lines are drawn:
The black line is the power. The light blue line is the turret room (temperature-
mean(temp))/1000 (so .001 step is 1 degree F).
Red. When the air conditioner console was turned on.
Green. When the compressor came on.
Dark blue. When the console and compressor were turned off.
Pink. When the 208 volt breaker for the air conditioner was turned off.
All of the plots show a step up in the power close to when the compressor
came on (red). All of the plots show a step down when the console/compressor
goes off (blue). The last 3 plots show a spike when the console is
turned off (before the drop). Turning off the 208 breaker makes no difference.
Plots 1,2,3 the louvre motors where turned on/off multiple times during
the 300 seconds without making a noticeable change in the power.
Plots 4,5 had the blower motor come on 60,30 seconds before the compressor
with no change in the power.
Plots 4,5 we delayed the compressor coming on by setting the temperature
threshold to 30 C. In plot 4 it took 70 seconds before the compressor
came on. In plot 5 it came on within 30 seconds.
The louvre motors and blower do not cause the jumps.
Turning the 208 breaker off makes no difference.
Turning the compressor/console off causes the power to jump back down.
On 05oct01 data the power jumps up when the compressor comes on. On the
28sep01 the power jumps up 10 seconds after the console comes on.
On 28sep01 the data was contiguous in time. After the compressor
was turned off, the unit would not allow the compressor to come back on
for 300 seconds (see the plot 28sep01 and the distance between the blue
line and the next green line).
On 05oct01 we did 5 minute position switching with
the telescope (to try and cancel ripples in the bandpass for the images).
After every 300 seconds of air conditioning on/off, there was a delay of
360 seconds where everything was off before the next cycle began.
This shows that the thing that makes the power jump
up must be some electronics that starts when the console wants to turn
the compressor on (e.g.. the temperature is above the turn on threshold).
The actual turning on of the compressor is delayed because of the safety
31oct01 Timer module for compressor causing the problem.
On 31oct01 blank sky was tracked using lbn. The correlator was setup
The Ac console sends 24 volts to the compressor to start. In the compressor
is an analog timer module that delays this 24 volts (for 5 minutes after
the last turn off so the pressures can equalize) before passing the 24
volts on to the contactor that starts the compressor.
sbc1 1432Mhz 12.5 Mhz bw
sbc2 1400Mhz 25.0 Mhz bw
sbc3 1413Mhz 12.5 Mhz bw
We inserted double pole, double throw switches before
and after the timer module. The datataking sequence was:
figure plots the total power versus time. The 3 colors are the 3 frequency
bands. The top plot is pol A while the bottom plot is pol B. The vertical
dashed lines bracket where the output switch from the timer module was
on. The total power jumps by about .005 Tsys (.005*32=.16Kelvin).
start taking data everything off.
@20 secs: turn on the power breaker.
@64 secs: switch on the ac console.
@175secs: switch on the input switch to the timer module.
@275secs: switch on the output switch from the timer module.
@384secs: switch off the output switch from the timer module.
@488secs: switch on the output switch from the timer module.
@532secs: switch off the output switch from the timer module.
@700secs: switch off the input switch to the timer module.
@841secs: turn off ac console.
The jump occurred when the output switch from the timer was thrown,
even though the 24 volts from the timer module was not being passed to
the compressor contactor (it hadn't timed out yet). The timer module
must be creating rfi that is being radiated out the wires that go to the
compressor. We completely bypassed the timer module and saw no jump in
the total power when the 24 volts was sent from the console to the compressor
contactor. The module has been removed and we are checking out how to "fix
it". The Ac can't be used until the timer is replaced.